OpenStudy (anonymous):

Find the number of sides for a regular polygon if the area is 480, the side length is 10 and the apothem is 6. Answer 30 16 8 4

OpenStudy (anonymous):

Area= (1/2)ap a=apothem p=perimeter We have: Area=480 a=6 we need to find p: \[480=(\frac{1}{2})(6)p \rightarrow \frac{480*2}{6}=p \rightarrow p=160\] since it is a regular polygon, all the sides have equal length, so 160=10x x=16 Thus the polygon has 16 sides